OpenStudy (anonymous):

Solve 4x2 = 12x + 40 Solving Quadratics by Factoring help please! :)

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20110714170233AAf81lE this might be the answer :P

OpenStudy (anonymous):

Take everything to one side 4x2 - 12x - 40 = 0 4x2 - 20x + 8x -40 = 0 4x(x-5) + 8(x-5) = 0 (4x+8)(x-5) = 0 x-5=0 x=5 4x+8 = 0 x = -2